# From Specific-MLLMs to Omni-MLLMs: A Survey on MLLMs Aligned with Multi-modalities

### v2

@@ Abstract
To tackle complex tasks in real-world scenarios, more researchers are focusing on Omni-MLLMs, which aim to achieve omni-modal understanding and generation. Beyond the constraints of any specific non-linguistic modality, Omni-MLLMs map various non-linguistic modalities into the embedding space of LLMs and enable the interaction and understanding of arbitrary combinations of modalities within a single model. In this paper, we systematically investigate relevant research and provide a comprehensive survey of Omni-MLLMs. Specifically, we first explain the four core components of Omni-MLLMs for unified multi-modal modeling with a meticulous taxonomy that offers novel perspectives. Then, we introduce the effective integration achieved through two-stage training and discuss the corresponding datasets as well as evaluation. Furthermore, we summarize the main challenges of current Omni-MLLMs and outline future directions. We hope this paper serves as an introduction for beginners and promotes the advancement of related research. Resources will be made public.
@@ 2 Omni-MLLM Architecture
Projection [ Cross-modal Understanding: eP-ALM (Shukor et al., 2023a),X-LLM (Chen et al., 2023a), ChatBridge
@@
 (Li et al., 2024i), DAMC (Chen et al., 2024a), CAT (Ye et al., 2024b), AVicuna (Tang et al., 2024a),
QaP (Liang et al., 2024), Uni-Moe (Li et al., 2024g), Video-LLaMA2 (Cheng et al., 2024b), Emotion-LLaMA
@@
 (Li et al., 2024h), VITA (Fu et al., 2024b), Mini-Omni2 (Xie and Wu, 2024), CAD-MLLM (Xu et al., 2024a),
@@
 (Luo et al., 2024), OccLLaMA (Wei et al., 2024), MIO (Wang et al., 2024g), Gesticulator (Pang et al., 2024),
@@ 2.2.1 Projection Alignment
It is also worth noting that in multi-branch Omni-MLLMs, different branches may utilize distinct implementations to better accommodate the unique characteristics of each modality (Li et al., 2024i). For example, Uni-MoE Li et al. (2024g) uses a linear projection for the visual modality and a Q-Former for the auditory modality. Meanwhile, uni-branch Omni-MLLMs, when using an attention-based projector, typically design multiple modality-specific learnable vectors to extract key information from various non-linguistic modalities (Yu et al., 2024a; Han et al., 2024a; Yu et al., 2024c).
@@ 2.3 Multi-modalities Interaction
For interaction, most Omni-MLLMs Chen et al. (2023a); Ye et al. (2024a); OpenGV (2024); Li et al. (2025) concatenate aligned non-linguistic modality features $\mathbf{F}_{p}$ with textual features $\mathbf{F}_{t}$ at the input level, enabling interaction in a progressive and layer-by-layer manner. Meanwhile, some works, such as ImageBind-LLM (Han et al., 2023) and TVL-LLaMA (Fu et al., 2024c), insert $\mathbf{F}_{p}$ into specific layers or all layers of the LLMs to mitigate the loss of modality information (Shukor et al., 2023a).
In terms of the number of modalities involved in interactions, compared to Specific-MLLMs that are limited to dual-modal interactions between a single non-linguistic modality and text (Liu et al., 2023c; Xu et al., 2024b), Omni-MLLMs not only support multiple dual-modal interactions but also enable omni-multimodal interactions involving more than two non-linguistic modalities (Zhao et al., 2023b; Wang et al., 2024g). For example, X-InstructBLIP (Panagopoulou et al., 2024) enables dual-modal interactions such as vision-text, audio-text, and 3D-text, as well as omni-modal interactions like vision-audio-text and 3D-vision-text, showcasing the ability of Omni-MLLMs to handle arbitrary combinations of modalities.
@@ 2.4 Multi-modalities Generation
This approach directly utilizes the discrete text output from the LLM to invoke Text-to-X generation models (Liu et al., 2023b; Luo et al., 2023c; Brooks et al., 2023) based on the content of the text. For example, VITA (Fu et al., 2024b) employs TTS tools (RVC-Boss, ) to convert the output text into corresponding speech, while ModelVerse (Wang et al., 2024c) and UnifiedMLLM (Li et al., 2024h) use the text to specify the generation model and utilize the corresponding descriptions to generate different modalities.
@@ 3.1.2 Output Alignment
The training of output alignment typically utilizes the same X-text paired dataset as input alignment and adheres to the identical training sequence. Meanwhile, the training objectives for output alignment vary depending on the multi-modality generation methods in Section 2.4. Token-based generative Omni-MLLMs optimize the extended LLM head by minimizing the text generation loss associated with modality-specific discrete tokens (Lu et al., 2024a; Wei et al., 2024). Representation-based generative Omni-MLLMs generally optimize their output projectors by minimizing the composite loss comprising three components (Xie and Wu, 2024; Yang et al., 2023b): 1) the text generation loss of signal tokens; 2) the L2 distance between the output representation and the condition vector of the corresponding decoder, i.e. MSE loss; and 3) the conditional latent denoising loss Rombach et al. (2022). For text-based generative Omni-MLLMs, as there is no additional output structure, the output alignment training is generally not required (Wang et al., 2024c; Li et al., 2024h).
@@ 3.2 Multi-modalities Instruction Fine-tuning
Compared to Specific-MLLMs, Omni-MLLMs not only leverage multiple uni-modal instruction data of different modalities for training but also use cross-modal instruction data to enhance their cross-modal ability (Ye et al., 2024a; Zhan et al., 2024; Li et al., 2024c). In addition to directly mixing different instruction data for training (Panagopoulou et al., 2024; Ye et al., 2024a), some works like Uni-Moe (Li et al., 2024g) and Lyra (Zhong et al., 2024) adopt a multi-step fine-tuning approach, introducing different uni-modal and cross-modal instruction data in a specific order for training to gradually enhance their uni-modal and cross-modal ability.
@@ 4.1 Training Data
Omni-MLLMs leverage caption datasets from various modalities to construct X-Text paired data for alignment pre-training, such as the WebVid (Bain et al., 2021) for visual modality and the AudioCaps (Kim et al., 2019) for auditory modality. However, for data-scarce modalities like depth maps and thermal maps, large-scale text-paired data is lacking (Zhu et al., 2024a; Girdhar et al., 2023). To address this, synthetic methods that use DPT models Ranftl et al. (2021); Bhat et al. (2023); Xu et al. (2023) or image translation models (Lee et al., 2023) to convert image-text pairs into other modality text pairs are widely employed Han et al. (2024a); Chen et al. (2024c); Jiang et al. (2024c).Moreover, interleaved datasets (Zhu et al., 2023a) are used for alignment pre-training in some works (Tang et al., 2024b) to enhance the contextual understanding capability.
@@
(1) Template-based Construction: Most works Sun et al. (2023a); Zhao et al. (2023b); Zhang et al. (2024b) utilize cross-modal downstream datasets (Sanabria et al., 2018; Chen et al., 2020b) combined with predefined templates to construct cross-modal instructions; (2) GPT Generation: Following the paradigm of LLaVA (Liu et al., 2023c), some Omni-MLLMs Lyu et al. (2023); Zhao et al. (2023b) leverage the labels from the annotated dataset (Lin et al., 2014; Bain et al., 2021) or use pre-trained models like SAM (Chen et al., 2023c) and GRIT (Wu et al., 2024a) to extract meta-information (e.g., captions and object categories) of different modalities. Then they employ powerful LLMs (OpenAI, 2023b, a) to generate cross-modal instructions based on the obtained meta-information; (3) T2X Generation: Li et al. (2024g) use TTS tools to convert the Image-Text2Text uni-modal instructions from LLaVA-v1.5 (Liu et al., 2024a) into Image-Speech-Text2Text cross-modal instructions. AnyGPT (Zhan et al., 2024) and NextGPT (Wu et al., 2024b) leverage Text2X models such as DALL-E-3 (Shi et al., 2020) and MusicGen (Copet et al., 2023) to convert the GPT-generated pure text instructions into Xs2Xs cross-modal instructions. Details about training data are shown in Appendix C.1
@@ 4.2 Benchmark
Uni-modal understanding assesses the ability of Omni-MLLMs to comprehend and reason on different non-linguistic modalities, including downstream X-Text2Text datasets such as X-Caption Plummer et al. (2015); Xu et al. (2016), X-QA Goyal et al. (2017); Xu et al. (2017a), and X-Classification Deitke et al. (2023), as well as comprehensive multi-task benchmarks Liu et al. (2024d); Fu et al. (2024a, 2023).
@@
Cross-modal understanding evaluates the ability of Omni-MLLMs to jointly comprehend and reason across multiple non-linguistic modalities like Image-Speech-Text2Text (Li et al., 2024g; OpenGV, 2024), Video-Audio-Text2Text (Li et al., 2022a, b), and Image-3D-Text2Text (Panagopoulou et al., 2024).
@@ 5.1 Expansion of modalities
Expanding new modalities may adjust the shared parameters, potentially causing catastrophic forgetting of previously trained modalities knowledge (Yu et al., 2024a). This issue can be partially mitigated by mixing trained modality data (Han et al., 2024a; Li et al., 2024g) or fine-tuning only the modality-specific parameters Yu et al. (2024a, c), but both approaches make the training process more complex.
@@
Although the data synthesis method in Section 4.1 can help alleviate the lack of text-paired data and instruction data for low-resource modalities (Han et al., 2024a; Jiang et al., 2024c), the absence of real modality may lead to biases in understanding of that modality.
@@ 5.2 Cross-modal capabilities
Although Omni-MLLMs employ various methods in Section 4.1 to generate cross-modal instruction data, there is still significant room for improvement and expansion, including enhancing the diversity of instructions, incorporating longer contextual dialogues, and exploring more diverse modality interaction paradigms. Similarly, cross-modal benchmarks such as OmniBench (Li et al., 2024f) and OmniR (Chen et al., 2024d) still fall short in terms of task richness and instruction diversity when compared to uni-modal benchmarks like MMMU (Yue et al., 2024) and MME (Fu et al., 2023). And the variety of modalities they cover is also relatively limited.
@@ Appendix A Related Survey
With the advent of MLLMs, there are several surveys detailing the current progress of MLLMs. Yin et al. (2023a); Wu et al. (2023); Caffagni et al. (2024) focus on the early Vision-MLLMs, while Çoban et al. (2024) and Ma et al. (2024) respectively summarize the Audio-MLLMs and 3D-MLLMs. Zhang et al. (2024a); Wang et al. (2024a) conduct an investigation into various Specific-MLLMs of different modalities. He et al. (2024); Chen et al. (2024b) discuss the expansion of MLLM’s generative capabilities. Some works discuss MLLMs in specific domains, such as medicine (Xiao et al., 2024), agriculture (Zhu et al., 2024c), and autonomous driving (Cui et al., 2024). Some works highlight some specific tasks such as safety (Fan et al., 2024), hallucination (Bai et al., 2024b), and acceleration (Zhu et al., 2024c). And Li and Lu (2024); Huang and Zhang (2024) focus on the evaluation of MLLM performance.
@@ B.1 Modality Encoder
AST (Gong et al., 2021), Beats (Chen et al., 2023d), Whisper (Radford et al., 2023), HuBert (Hsu et al., 2021), CLAP (Elizalde et al., 2023), Conformer (Gulati et al., 2020), MERT (Li et al., 2024e), and PANN (Kong et al., 2020) encode the audio modality to obtain continuous features.
@@ C.1 Details of Training Data
LAION-115M (Schuhmann et al., 2021a) X-Text Image,Text 115M
@@
GroundingGPT’s SFT (Li et al., 2024i) Flickr30K, VCR, Activitynet Captions, Clotho… Uni-Modal Understanding Image,Video,Audio,Text GPT Instructionalization 1M
@@
Uni-Moe’s SFT (Li et al., 2024g) LLaVA-Instruct-150K, LibriSpeech, VideoInstruct100K… Uni-Modal Understanding,Cross-Modal Understanding Image,Video,Audio,Text Template Instructionalization, T2X generation 874K
@@
UnifiedMLLM’s SFT (Li et al., 2024h) LISA,SmartEdit… Uni-Modal Understanding,Cross-Modal Understanding, Uni-Modal Generation,Cross-Modal Generation Image,Video,Audio,Text Template Instructionalization, GPT Generation 100K
@@ C.2 Details of Benchmark
POPE (Li et al., 2023b) Unimodal Understanding Image,Text Hallucination Acc
@@
MSVD-QA (Xu et al., 2017a) Unimodal Understanding Video,Text QA Acc
@@
MSVD-Caption (Xu et al., 2017a) Unimodal Understanding Video,Text QA CIDER
@@
EHSL (Li et al., 2024g) Unimodal Understanding Audio,Text QA Acc
@@
RACE-Audio (Li et al., 2024g) Crossmodal Understanding Image,Audio,Text Comprehensive Benchmark Acc
@@
MMBench-Audio (Li et al., 2024g) Crossmodal Understanding Image,Audio,Text Comprehensive Benchmark Acc

### v3

@@ Abstract
To tackle complex tasks in real-world scenarios, more researchers are focusing on Omni-MLLMs, which aim to achieve omni-modal understanding and generation. Beyond the constraints of any specific non-linguistic modality, Omni-MLLMs map various non-linguistic modalities into the embedding space of LLMs and enable the interaction and understanding of arbitrary combinations of modalities within a single model. In this paper, we systematically investigate relevant research and provide a comprehensive survey of Omni-MLLMs. Specifically, we first explain the four core components of Omni-MLLMs for unified multi-modal modeling with a meticulous taxonomy that offers novel perspectives. Then, we introduce the effective integration achieved through two-stage training and discuss the corresponding datasets as well as evaluation. Furthermore, we summarize the main challenges of current Omni-MLLMs and outline future directions. We hope this paper serves as an introduction for beginners and promotes the advancement of related research. Resources have been made publicly available at https://github.com/threegold116/Awesome-Omni-MLLMs.
@@ 2 Omni-MLLM Architecture
Projection [ Cross-modal Understanding: eP-ALM (Shukor et al., 2023),X-LLM (Chen et al., 2023a), ChatBridge
@@
 (Li et al., 2024h), DAMC (Chen et al., 2024a), CAT (Ye et al., 2024b), AVicuna (Tang et al., 2024a),
QaP (Liang et al., 2024), Uni-Moe (Li et al., 2024f), Video-LLaMA2 (Cheng et al., 2024b), Emotion-LLaMA
@@
 (Li et al., 2024g), VITA (Fu et al., 2024b), Mini-Omni2 (Xie and Wu, 2024), CAD-MLLM (Xu et al., 2024a),
@@
 (Luo et al., 2024), OccLLaMA (Wei et al., 2024), MIO (Wang et al., 2024f), Gesticulator (Pang et al., 2024),
@@ 2.2.1 Projection Alignment
It is also worth noting that in multi-branch Omni-MLLMs, different branches may utilize distinct implementations to better accommodate the unique characteristics of each modality (Li et al., 2024h). For example, Uni-MoE Li et al. (2024f) uses a linear projection for the visual modality and a Q-Former for the auditory modality. Meanwhile, uni-branch Omni-MLLMs, when using an attention-based projector, typically design multiple modality-specific learnable vectors to extract key information from various non-linguistic modalities (Yu et al., 2024a; Han et al., 2024a; Yu et al., 2024c).
@@ 2.3 Multi-modalities Interaction
For interaction, most Omni-MLLMs Chen et al. (2023a); Ye et al. (2024a); OpenGV (2024); Li et al. (2025) concatenate aligned non-linguistic modality features $\mathbf{F}_{p}$ with textual features $\mathbf{F}_{t}$ at the input level, enabling interaction in a progressive and layer-by-layer manner. Meanwhile, some works, such as ImageBind-LLM (Han et al., 2023) and TVL-LLaMA (Fu et al., 2024c), insert $\mathbf{F}_{p}$ into specific layers or all layers of the LLMs to mitigate the loss of modality information (Shukor et al., 2023).
In terms of the number of modalities involved in interactions, compared to Specific-MLLMs that are limited to dual-modal interactions between a single non-linguistic modality and text (Liu et al., 2023c; Xu et al., 2024b), Omni-MLLMs not only support multiple dual-modal interactions but also enable omni-multimodal interactions involving more than two non-linguistic modalities (Zhao et al., 2023b; Wang et al., 2024f). For example, X-InstructBLIP (Panagopoulou et al., 2024) enables dual-modal interactions such as vision-text, audio-text, and 3D-text, as well as omni-modal interactions like vision-audio-text and 3D-vision-text, showcasing the ability of Omni-MLLMs to handle arbitrary combinations of modalities.
@@ 2.4 Multi-modalities Generation
This approach directly utilizes the discrete text output from the LLM to invoke Text-to-X generation models (Liu et al., 2023b; Luo et al., 2023c; Brooks et al., 2023) based on the content of the text. For example, VITA (Fu et al., 2024b) employs TTS tools (RVC-Boss, ) to convert the output text into corresponding speech, while ModelVerse (Wang et al., 2024c) and UnifiedMLLM (Li et al., 2024g) use the text to specify the generation model and utilize the corresponding descriptions to generate different modalities.
@@ 3.1.2 Output Alignment
The training of output alignment typically utilizes the same X-text paired dataset as input alignment and adheres to the identical training sequence. Meanwhile, the training objectives for output alignment vary depending on the multi-modality generation methods in Section 2.4. Token-based generative Omni-MLLMs optimize the extended LLM head by minimizing the text generation loss associated with modality-specific discrete tokens (Lu et al., 2024a; Wei et al., 2024). Representation-based generative Omni-MLLMs generally optimize their output projectors by minimizing the composite loss comprising three components (Xie and Wu, 2024; Yang et al., 2023b): 1) the text generation loss of signal tokens; 2) the L2 distance between the output representation and the condition vector of the corresponding decoder, i.e. MSE loss; and 3) the conditional latent denoising loss Rombach et al. (2022). For text-based generative Omni-MLLMs, as there is no additional output structure, the output alignment training is generally not required (Wang et al., 2024c; Li et al., 2024g).
@@ 3.2 Multi-modalities Instruction Fine-tuning
Compared to Specific-MLLMs, Omni-MLLMs not only leverage multiple uni-modal instruction data of different modalities for training but also use cross-modal instruction data to enhance their cross-modal ability (Ye et al., 2024a; Zhan et al., 2024; Li et al., 2024c). In addition to directly mixing different instruction data for training (Panagopoulou et al., 2024; Ye et al., 2024a), some works like Uni-Moe (Li et al., 2024f) and Lyra (Zhong et al., 2024) adopt a multi-step fine-tuning approach, introducing different uni-modal and cross-modal instruction data in a specific order for training to gradually enhance their uni-modal and cross-modal ability.
@@ 4.1 Training Data
Omni-MLLMs leverage caption datasets from various modalities to construct X-Text paired data for alignment pre-training, such as the WebVid (Bain et al., 2021) for visual modality and the AudioCaps (Kim et al., 2019) for auditory modality. However, for data-scarce modalities like depth maps and thermal maps, large-scale text-paired data is lacking (Zhu et al., 2024a; Girdhar et al., 2023). To address this, synthetic methods that use DPT models Ranftl et al. (2021); Bhat et al. (2023); Xu et al. (2023) or image translation models (Lee et al., 2023) to convert image-text pairs into other modality text pairs are widely employed Han et al. (2024a); Chen et al. (2024c); Zhu et al. (2024a). Moreover, interleaved datasets (Zhu et al., 2023a) are used for alignment pre-training in some works (Tang et al., 2024b) to enhance the contextual understanding capability.
@@
(1) Template-based Construction: Most works Sun et al. (2023a); Zhao et al. (2023b); Zhang et al. (2024b) utilize cross-modal downstream datasets (Sanabria et al., 2018; Chen et al., 2020b) combined with predefined templates to construct cross-modal instructions; (2) GPT Generation: Following the paradigm of LLaVA (Liu et al., 2023c), some Omni-MLLMs Lyu et al. (2023); Zhao et al. (2023b) leverage the labels from the annotated dataset (Lin et al., 2014; Bain et al., 2021) or use pre-trained models like SAM (Chen et al., 2023c) and GRIT (Wu et al., 2024a) to extract meta-information (e.g., captions and object categories) of different modalities. Then they employ powerful LLMs (OpenAI, 2023b, a) to generate cross-modal instructions based on the obtained meta-information; (3) T2X Generation: Li et al. (2024f) use TTS tools to convert the Image-Text2Text uni-modal instructions from LLaVA-v1.5 (Liu et al., 2024a) into Image-Speech-Text2Text cross-modal instructions. AnyGPT (Zhan et al., 2024) and NextGPT (Wu et al., 2024b) leverage Text2X models such as DALL-E-3 (Shi et al., 2020) and MusicGen (Copet et al., 2023) to convert the GPT-generated pure text instructions into Xs2Xs cross-modal instructions. Details about training data are shown in Appendix C.1
@@ 4.2 Benchmark
Uni-modal understanding assesses the ability of Omni-MLLMs to comprehend and reason on different non-linguistic modalities, including downstream X-Text2Text datasets such as X-Caption Plummer et al. (2015); Xu et al. (2016), X-QA Goyal et al. (2017); Xu et al. (2017), and X-Classification Deitke et al. (2023), as well as comprehensive multi-task benchmarks Liu et al. (2024d); Fu et al. (2024a, 2023).
@@
Cross-modal understanding evaluates the ability of Omni-MLLMs to jointly comprehend and reason across multiple non-linguistic modalities like Image-Speech-Text2Text (Li et al., 2024f; OpenGV, 2024), Video-Audio-Text2Text (Li et al., 2022a, b), and Image-3D-Text2Text (Panagopoulou et al., 2024).
@@ 5.1 Expansion of modalities
Expanding new modalities may adjust the shared parameters, potentially causing catastrophic forgetting of previously trained modalities knowledge (Yu et al., 2024a). This issue can be partially mitigated by mixing trained modality data (Han et al., 2024a; Li et al., 2024f) or fine-tuning only the modality-specific parameters Yu et al. (2024a, c), but both approaches make the training process more complex.
@@
Although the data synthesis method in Section 4.1 can help alleviate the lack of text-paired data and instruction data for low-resource modalities (Han et al., 2024a; infraredllava), the absence of real modality may lead to biases in understanding of that modality.
@@ 5.2 Cross-modal capabilities
Although Omni-MLLMs employ various methods in Section 4.1 to generate cross-modal instruction data, there is still significant room for improvement and expansion, including enhancing the diversity of instructions, incorporating longer contextual dialogues, and exploring more diverse modality interaction paradigms. Similarly, cross-modal benchmarks such as OmniBench (Li et al., 2024e) and OmniR (Chen et al., 2024d) still fall short in terms of task richness and instruction diversity when compared to uni-modal benchmarks like MMMU (Yue et al., 2024) and MME (Fu et al., 2023). And the variety of modalities they cover is also relatively limited.
@@ Appendix A Related Survey
With the advent of MLLMs, there are several surveys detailing the current progress of MLLMs. Yin et al. (2023a); Wu et al. (2023); Caffagni et al. (2024) focus on the early Vision-MLLMs, while Çoban et al. (2024) and Ma et al. (2024) respectively summarize the Audio-MLLMs and 3D-MLLMs. Zhang et al. (2024a); Wang et al. (2024a) conduct an investigation into various Specific-MLLMs of different modalities. He et al. (2024); Chen et al. (2024b) discuss the expansion of MLLM’s generative capabilities. Some works discuss MLLMs in specific domains, such as medicine (Xiao et al., 2024), agriculture (Zhu et al., 2024b), and autonomous driving (Cui et al., 2024). Some works highlight some specific tasks such as safety (Fan et al., 2024), hallucination (Bai et al., 2024b), and acceleration (Zhu et al., 2024b). And Li and Lu (2024); Huang and Zhang (2024) focus on the evaluation of MLLM performance.
@@ B.1 Modality Encoder
AST (Gong et al., 2021), Beats (Chen et al., 2023d), Whisper (Radford et al., 2023), HuBert (Hsu et al., 2021), CLAP (Elizalde et al., 2023), Conformer (Gulati et al., 2020), MERT (Li et al., 2024d), and PANN (Kong et al., 2020) encode the audio modality to obtain continuous features.
@@ C.1 Details of Training Data
LAION-115M (Schuhmann et al., 2021) X-Text Image,Text 115M
@@
GroundingGPT’s SFT (Li et al., 2024h) Flickr30K, VCR, Activitynet Captions, Clotho… Uni-Modal Understanding Image,Video,Audio,Text GPT Instructionalization 1M
@@
Uni-Moe’s SFT (Li et al., 2024f) LLaVA-Instruct-150K, LibriSpeech, VideoInstruct100K… Uni-Modal Understanding,Cross-Modal Understanding Image,Video,Audio,Text Template Instructionalization, T2X generation 874K
@@
UnifiedMLLM’s SFT (Li et al., 2024g) LISA,SmartEdit… Uni-Modal Understanding,Cross-Modal Understanding, Uni-Modal Generation,Cross-Modal Generation Image,Video,Audio,Text Template Instructionalization, GPT Generation 100K
@@ C.2 Details of Benchmark
POPE (Li et al., 2023) Unimodal Understanding Image,Text Hallucination Acc
@@
MSVD-QA (Xu et al., 2017) Unimodal Understanding Video,Text QA Acc
@@
MSVD-Caption (Xu et al., 2017) Unimodal Understanding Video,Text QA CIDER
@@
EHSL (Li et al., 2024f) Unimodal Understanding Audio,Text QA Acc
@@
RACE-Audio (Li et al., 2024f) Crossmodal Understanding Image,Audio,Text Comprehensive Benchmark Acc
@@
MMBench-Audio (Li et al., 2024f) Crossmodal Understanding Image,Audio,Text Comprehensive Benchmark Acc